Evaluate the integral. (Hint: Substitution Rule] ſ(27 (2.c + 3) (2.2 +6x + 1)dx
Answer #1

Solu given 5 (2x+3) (2x2 +62+1) & dex het =t 2x2 +6X+1 (42+6) dx =de - =dd =2(2x+3)dz => (2x+3) de - so our Ź dt integration